# Data-Driven Fairness Generalization for Deepfake Detection

### v1

@@ 2.1 Deepfake Detection
To train deepfake detection models that perform well during training and in practice; on unseen datasets, several studies [Li et al., 2020b, Zhao et al., 2021b, Guan et al., 2022, Li, 2018] have introduced novel methods for manually synthesizing a variety of face forgeries similar to deepfakes. These methods help deepfake detection models learn more generalized representations of artifacts, improving their ability to identify deepfakes across different scenarios.
The authors in [Chen et al., 2022] proposed enhancing the diversity of forgeries through adversarial training to improve the robustness of deepfake detectors in recognizing various forgeries. FaceCutout [Das et al., 2021] uses facial landmarks and randomly cuts out different parts of the face (such as the mouth, eye, etc.) to improve the robustness of deepfake detection. Face-Xray [Li et al., 2020b] on the other hand, involves generating blended images (BI) by combining two different faces using a global transformation and then training a model to distinguish between real and blended faces. The authors in [Shiohara and Yamasaki, 2022] expanded on this concept by creating a synthetic training dataset with self-blended images (SBI), which are created by applying a series of data augmentation techniques to real images. SBI has demonstrated even better generalization to previously unseen deepfakes compared to Face-Xray [Li et al., 2020b]. After synthesizing forged images (i.e., pseudo deepfakes) using specialized augmentation techniques, they trained a binary classification model to perform the detection task.
@@ 4.1 Experimental Settings
To evaluate the fairness generalization capability of our proposed approach, we train our model on the widely used FaceForensics++ (FF++) dataset [Rossler et al., 2019]. For testing, we use FF++, Deepfake Detection Challenge (DFDC) [Dolhansky et al., 2020], and Celeb-DF [Li et al., 2020c]. We used the real images from these datasets and created synthetic images as the fake images using the method in Section 3.2.
Since the original datasets do not include demographic information, we follow the approach in [Ju et al., 2024] for data processing and annotation. The datasets are categorized by race and gender, forming the following intersection groups: Black-Male (B-M), Black-Female (B-F), White-Male (W-M), White-Female (W-F), Asian-Male (A-M), Asian-Female (A-F), Other-Male (O-M), and Other-Female (O-F). The Celeb-DF [Li et al., 2020c] does not contain the Asian sub-group.
@@ 4.2.1 Cross-dataset Evaluation Results
DFDC [Dolhansky et al., 2020] Celeb-DF [Li et al., 2020c]
@@
Table 2: Cross-dataset evaluation across different methods for deepfake detection. The methods were trained on (FF++ [Rossler et al., 2019]) and tested on DFDC [Dolhansky et al., 2020] and Celeb-DF [Li et al., 2020c].
@@
Figure 8: Cross-dataset evaluation across different methods. Methods are trained on (FF++ [Rossler et al., 2019]) and tested on Celeb-DF [Li et al., 2020c]. Performance comparison is across gender. See Table 2 for exact numerical values
@@
Figure 9: Cross-dataset evaluation across different methods. Methods are trained on (FF++ [Rossler et al., 2019]) and tested on Celeb-DF [Li et al., 2020c]. Performance comparison is across races. See Table 2 for exact numerical values

### v2

@@ 2.1 Deepfake Detection
To train deepfake detection models that perform well during training and in practice; on unseen datasets, several studies [Li et al., 2020a, Zhao et al., 2021b, Guan et al., 2022, Li, 2018] have introduced novel methods for manually synthesizing a variety of face forgeries similar to deepfakes. These methods help deepfake detection models learn more generalized representations of artifacts, improving their ability to identify deepfakes across different scenarios.
The authors in [Chen et al., 2022] proposed enhancing the diversity of forgeries through adversarial training to improve the robustness of deepfake detectors in recognizing various forgeries. FaceCutout [Das et al., 2021] uses facial landmarks and randomly cuts out different parts of the face (such as the mouth, eye, etc.) to improve the robustness of deepfake detection. Face-Xray [Li et al., 2020a] on the other hand, involves generating blended images (BI) by combining two different faces using a global transformation and then training a model to distinguish between real and blended faces. The authors in [Shiohara and Yamasaki, 2022] expanded on this concept by creating a synthetic training dataset with self-blended images (SBI), which are created by applying a series of data augmentation techniques to real images. SBI has demonstrated even better generalization to previously unseen deepfakes compared to Face-Xray [Li et al., 2020a]. After synthesizing forged images (i.e., pseudo deepfakes) using specialized augmentation techniques, they trained a binary classification model to perform the detection task.
@@ 4.1 Experimental Settings
To evaluate the fairness generalization capability of our proposed approach, we train our model on the widely used FaceForensics++ (FF++) dataset [Rossler et al., 2019]. For testing, we use FF++, Deepfake Detection Challenge (DFDC) [Dolhansky et al., 2020], and Celeb-DF [Li et al., 2020b]. We used the real images from these datasets and created synthetic images as the fake images using the method in Section 3.2.
Since the original datasets do not include demographic information, we follow the approach in [Ju et al., 2024] for data processing and annotation. The datasets are categorized by race and gender, forming the following intersection groups: Black-Male (B-M), Black-Female (B-F), White-Male (W-M), White-Female (W-F), Asian-Male (A-M), Asian-Female (A-F), Other-Male (O-M), and Other-Female (O-F). The Celeb-DF [Li et al., 2020b] does not contain the Asian sub-group.
@@ 4.2.1 Cross-dataset Evaluation Results
DFDC [Dolhansky et al., 2020] Celeb-DF [Li et al., 2020b]
@@
Table 2: Cross-dataset evaluation across different methods for deepfake detection. The methods were trained on (FF++ [Rossler et al., 2019]) and tested on DFDC [Dolhansky et al., 2020] and Celeb-DF [Li et al., 2020b].
@@
Figure 8: Cross-dataset evaluation across different methods. Methods are trained on (FF++ [Rossler et al., 2019]) and tested on Celeb-DF [Li et al., 2020b]. Performance comparison is across gender. See Table 2 for exact numerical values
@@
Figure 9: Cross-dataset evaluation across different methods. Methods are trained on (FF++ [Rossler et al., 2019]) and tested on Celeb-DF [Li et al., 2020b]. Performance comparison is across races. See Table 2 for exact numerical values